# Multiple peaks in gravitational waves induced from primordial curvature perturbations with non-Gaussianity

### v1

@@ 1 Introduction
Gravitational waves (GWs) can serve as a powerful probe into the fundamental physics in the early universe, like the inflation [1, 2], phase transitions [3, 4, 5], topological defects [5, 6], reheating/preheating era [7, 8], to name just a few. As there will be many more GW detectors in the future, such as LISA [9], Taiji [10, 11], Tianqin [12], ET [13], DECIGO [14, 15], and AION/MAGIS [16], it is, therefore, crucial to distinguish all different kinds of possible sources of stochastic GW backgrounds [17, 18]. For instance, the anisotropies in the stochastic GW backgrounds can serve as a characteristic probe among different sources of stochastic GW backgrounds [19, 20, 21].
One of these intriguing GW sources is the scalar-induced GWs [22, 23] (see [24] for a review) from having a sizeable enhancement in the power spectrum of primordial curvature perturbations at small scales in addition to the nearly scale-invariant one at large scales [25, 26] as measured from the cosmic microwave background (CMB) by the Planck team. Such a sizeable enhancement can also lead to productions of primordial black holes (PBHs), which might constitute all of or part of the dark matter components [27, 28, 29, 30, 31] but with the abundance sensitive externally to the choice of window functions [32] as well as the reheating history normalization [33] so long as the inflationary scale has not been fixed yet.
If there is indeed such an enhancement at small scales, then it is essential to recover the small-scale structures in the primordial curvature perturbations from the observed characteristics in the scalar-induced second-order GWs, and vice versa. A widely studied feature is the imprint from non-Gaussianity [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]  [^2] See, however, Refs. [46, 47, 48, 45] for the non-Gaussianity in the induced GWs. and anisotropies [19, 49, 50, 51, 52]. Higher-order calculations, like the third-order results [53, 54, 55], have found that, for primordial power spectrum with a delta peak, there are two peaks in the GW spectrum to the third order instead of a single peak to the second order. Recent works on the missing one-loop contributions in second-order induced GWs [56, 57, 58, 59, 60, 61, 62] may lead to a scale-invariant negative contribution in the infrared region in addition to the universal infrared behavior [63].
@@ 2 Scalar-induced gravitational waves
In this section, we start by reviewing the general formalism for the SIGW in the Gaussian case [71]. Later in this section, we turn to the case with non-Gaussinanity [38].
@@ 2.1 SIGW in Gaussian case
$\langle h_{\lambda_{1}}(\tau,\mathbf{k})h_{\lambda_{2}}(\tau,\mathbf{q})\rangle=\delta^{3}(\mathbf{k}+\mathbf{q})\delta^{\lambda_{1}\lambda_{2}}\frac{2\pi^{2}}{k^{3}}\Delta_{\lambda_{1}}(\tau,k).$ (2.5)
@@
It is conventional to introduce two new variables $u=|k-q|/k$ and $v=q/k$ to simplify the above equations. Then, combining (2.6), (2.15), (2.18), and (2.20)) and considering the limit $x\rightarrow\infty$ as we are interested in the GW spectrum observed today, we finally obtain [71, 38, 51]
@@ 2.2 SIGW in non-Gaussian case
In the non-Gaussian case, the simple splitting (2.20) does not hold anymore, and we consider a local-type non-Gaussianity  [^4] In some works, $f_{\mathrm{NL}}$ is used, which is related to ours by $F_{\mathrm{NL}}=\frac{3}{5}f_{\mathrm{NL}}$. up to the second order,
@@
×$\qquad\times v_{1}^{2}u_{1}^{2}\frac{\Delta_{g}^{2}(v_{1}v_{2}k)}{(v_{1}v_{2})^{2}}\frac{\Delta_{g}^{2}(v_{1}u_{2}k)}{(v_{1}u_{2})^{2}}\frac{\Delta_{g}^{2}(u_{1}v_{3}k)}{(u_{1}v_{3})^{2}}\frac{\Delta_{g}^{2}(u_{1}u_{3}k)}{(u_{1}u_{3})^{2}}.$ (2.38)
For the connected terms, we first introduce another three sets of variables $(u_{i},v_{i})$,
@@ 3.1 Multiple peaks in Gaussian case
$J(t_{1},s_{1},t_{2},s_{2})=\overline{J^{2}\left(\frac{t_{1}+s_{1}+1}{2},\frac{t_{1}-s_{1}+1}{2},\frac{t_{2}+s_{2}+1}{2},\frac{t_{2}-s_{2}+1}{2},x\rightarrow\infty\right)}.$ (3.3)
From the definition of $J(t_{1},s_{1},t_{2},s_{2})$, it is easy to see that it has poles in $t_{1,2}=\sqrt{3}-1$, therefore, the peak will appear at $\frac{1}{\tilde{k}_{i}}+\frac{1}{\tilde{k}_{j}}=\sqrt{3}$, and the step function further constrains the domain of $k$ in the range $|k_{*i}-k_{*j}|<k<k_{*i}+k_{*j}$. As a result, we will have at most $C_{n}^{2}$ and at least $n$ peaks. To be more specific, let $k_{*j}=nk_{*i}$ and $n>1$, then all the possible additional peaks are located at $k=\frac{n+1}{\sqrt{3}}k_{*i}<(n+1)k_{*i}$, and in order to have more peaks, it requires $\frac{n+1}{\sqrt{3}}k_{*i}>(n-1)k_{*i}$, namely,
@@ 3.2 Multiple peaks in non-Gaussian case
Due to the multiple integrals, it becomes more complicated for the non-Gaussian case. For $\Omega_{\mathrm{GW}}^{\mathrm{H}}$, after inserting (3.1) into (2.37), we can obtain
@@
×$\times\Theta\left(\frac{1}{\tilde{k}_{i}}+\frac{1}{\tilde{k}_{j}}+\frac{1}{\tilde{k}_{l}}-t-1\right)\Theta\left(t+1-\frac{1}{\tilde{k}_{j}}-\left|\frac{1}{\tilde{k}_{l}}-\frac{1}{\tilde{k}_{i}}\right|\right)\Theta\left(1-\left|\frac{2}{\tilde{k}_{j}}-t-1\right|\right),$
and for later convenience, we further define
@@
Since the integral over a logarithmic divergence can still be finite, we cannot simply trust the poles of $J(t_{1},s_{1},t_{2},s_{2})$ as the peak locations. However, based on following two observations: (i) the integral over a logarithmic divergence is $\mathcal{O}(1)$ and the peak of $\Omega_{\mathrm{GW}}^{\mathrm{H}}$ is also $\mathcal{O}(1)$ when $A=F_{\mathrm{NL}}=1$; (ii) in terms of analyzing the zero point of $J(t_{1},s_{1},t_{2},s_{2})$, we find the upper bound of the integral variable $t$ is always smaller than the biggest zero point (see Appendix A for details), that is to say, including the logarithmic divergence will significantly enhance the integral result, otherwise, the integral result is very small; therefore, we can see that the sufficient condition for having a peak is to include the logarithmic divergent pole into the domain of $t$, and hence we can obtain the domain of peaks as shown shortly below.
@@
This is very different from the Gaussian case, for example, when $m=1$, it roughly gives $n<7.46$, that is, when $\Omega_{\mathrm{GW}}^{\mathrm{G}}$ only has two peaks, $\Omega_{\mathrm{GW}}^{\mathrm{H}}$ can still have more than two peaks  [^6] In fact, this is a slightly weak condition, as it requires roughly $n<6$ to give a sizeable peak.. In the left panel of Fig. 2, we present an example of induced GWs from primordial curvature perturbations with a double-$\delta$ peak. Taking $k_{*2}=4k_{*1}$, for instance, $\Omega_{\mathrm{GW}}^{\mathrm{G}}$ admits only two peaks while $\Omega_{\mathrm{GW}}^{\mathrm{H}}$ has at least three peaks (the other two peaks are not obvious). In the right panel of Fig. 2, we let $m=1$ and vary $n$ from 7.5 to 6 to show the peak behavior.
Due to the existence of integral, the peak of $\Omega_{\mathrm{GW}}^{\mathrm{H}}$ is usually broader than $\Omega_{\mathrm{GW}}^{\mathrm{G}}$, and hence it would be more practical to narrow down a tighter domain of peaks for $\Omega_{\mathrm{GW}}^{\mathrm{H}}$ rather than finding the precise locations of peaks as it is much more difficult. After some efforts, we can obtain such domains of peaks (see the Appendix A for details) as

### v2

@@ 1 Introduction
Gravitational waves (GWs) can serve as a powerful probe into the fundamental physics in the early universe, such as inflation [1, 2], phase transitions [3, 4, 5], topological defects [5, 6], reheating/preheating era [7, 8], to name just a few. As there will be many more GW detectors in the future, such as LISA [9], Taiji [10, 11], Tianqin [12], ET [13], DECIGO [14, 15], and AION/MAGIS [16], it is, therefore, crucial to distinguish all different kinds of possible sources of stochastic GW backgrounds [17, 18]. For instance, the anisotropies in the stochastic GW backgrounds can serve as a characteristic probe among different sources of stochastic GW backgrounds [19, 20, 21].
The cosmic microwave background (CMB) measurement by the Planck team [22, 23] has preferred an adiabatic Gaussian random perturbation field for the primordial curvature perturbations with a nearly scale-invariant power spectrum at large scales, but the small upper bound for the tensor-to-scalar ratio has indicated a negligible amount of primordial GW background at the first order. However, a scalar-induced GW background [24, 25] (see [26] for a review) can be generated at the second order from the primordial curvature perturbations if there is a sizeable enhancement in the power spectrum of primordial curvature perturbations at small scales. Such a sizeable enhancement can also lead to productions of primordial black holes (PBHs), which could constitute part of, if not all, the dark matter components [27, 28, 29, 30, 31]. Note that the abundance of PBHs is sensitive externally to the choice of window functions [32] as well as the reheating history normalization [33] so long as the inflationary scale has not been fixed yet.
If there is indeed such an enhancement at small scales, then it is essential to recover the small-scale structures in the primordial curvature perturbations from the observed characteristics in the scalar-induced second-order GWs, and vice versa. A widely studied feature is the imprint from non-Gaussianity [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]  [^2] See, however, Refs. [47, 48, 49, 45] for the non-Gaussianity in the induced GWs. and anisotropies [19, 50, 51, 52, 53]. Higher-order calculations, like the third-order results [54, 55, 56], have found that, for primordial power spectrum with a delta peak, there are two peaks in the GW spectrum to the third order instead of a single peak to the second order. Existence of ultra-light PBHs ($M_{\mathrm{PBH}}<10^{9}g$) can also induces double peaks in GW energy spectrum [57]. Recent works on the missing one-loop contributions in second-order induced GWs [58, 59, 60, 61, 62, 63, 64] may lead to a scale-invariant negative contribution in the infrared region in addition to the universal infrared behavior [65].
@@ 2 Scalar-induced gravitational waves
In this section, we start by reviewing the general formalism for the SIGW in the Gaussian case [73, 74]. Later in this section, we turn to the case with non-Gaussianity following the treatment of Ref. [38] but with multiple peaks for the power spectrum of primordial curvature perturbation.
@@ 2.1 SIGW in Gaussian case
$\langle h_{\lambda_{1}}(\tau,\mathbf{k})h_{\lambda_{2}}(\tau,\mathbf{q})\rangle=\delta^{3}(\mathbf{k}+\mathbf{q})\delta^{\lambda_{1}\lambda_{2}}\frac{2\pi^{2}}{k^{3}}\Delta_{\lambda_{1}}^{2}(\tau,k).$ (2.5)
@@
It is conventional to introduce two new variables $u=|k-q|/k$ and $v=q/k$ to simplify the above equations. Then, combining (2.6), (2.15), (2.18), and (2.20) and considering the limit $x\rightarrow\infty$ as we are interested in the GW spectrum observed today, we finally obtain [73, 38, 52]
@@ 2.2 SIGW in non-Gaussian case
In the non-Gaussian case, the simple splitting (2.20) does not hold anymore, and we consider a local-type non-Gaussianity  [^4] In some other works, $f_{\mathrm{NL}}$ is used, which is related to ours by $F_{\mathrm{NL}}=\frac{3}{5}f_{\mathrm{NL}}$. For those who are interested in other types of non-Gaussianity, please refer to Refs. [46, 48]. up to the second order,
@@
×$\qquad\times v_{1}^{2}u_{1}^{2}\frac{\Delta_{g}^{2}(v_{1}v_{2}k)}{(v_{1}v_{2})^{2}}\frac{\Delta_{g}^{2}(v_{1}u_{2}k)}{(v_{1}u_{2})^{2}}\frac{\Delta_{g}^{2}(u_{1}v_{3}k)}{(u_{1}v_{3})^{2}}\frac{\Delta_{g}^{2}(u_{1}u_{3}k)}{(u_{1}u_{3})^{2}},$ (2.38)
where $\overline{J^{2}(u_{1},v_{1},u_{2},v_{2},x\rightarrow\infty)}$ is defined in (2.22). For the connected terms, we first introduce another three sets of variables $(u_{i},v_{i})$,
@@ 3.1 Multiple peaks in Gaussian case
$J(t_{1},s_{1},t_{2},s_{2})=\overline{J^{2}\left(\frac{t_{1}+s_{1}+1}{2},\frac{t_{1}-s_{1}+1}{2},\frac{t_{2}+s_{2}+1}{2},\frac{t_{2}-s_{2}+1}{2},x\rightarrow\infty\right)},$ (3.3)
and what we need is $J(t,s,t,s)$ with $s\in[-1,1]$ and $t\geq 0$:
$+\left((s^{2}-(1+t)^{2})+\frac{(t^{2}+2t+s^{2}-5)}{4}\log\left|\frac{(t+1)^{2}-3}{(s^{2}-3)^{2}}\right|\right)^{2}\bigg{]}.$ (3.4)
From the definition of $J(t,s,t,s)$, it is easy to see that it has poles in $t=\sqrt{3}-1$, therefore, the peak will appear at $\frac{1}{\tilde{k}_{i}}+\frac{1}{\tilde{k}_{j}}=\sqrt{3}$, and the step function further constrains the domain of $k$ in the range $|k_{*i}-k_{*j}|<k<k_{*i}+k_{*j}$. As a result, we will have at most $C_{n}^{2}$ and at least $n$ peaks. To be more specific, let $k_{*j}=nk_{*i}$ and $n>1$, then all the possible additional peaks are located at $k=\frac{n+1}{\sqrt{3}}k_{*i}<(n+1)k_{*i}$, and in order to have more peaks, it requires $\frac{n+1}{\sqrt{3}}k_{*i}>(n-1)k_{*i}$, namely,
@@ 3.2 Multiple peaks in non-Gaussian case
Due to the increasing number of integral variables, the calculation of $\Omega_{\mathrm{GW}}$ becomes more complicated in the non-Gaussian case. For $\Omega_{\mathrm{GW}}^{\mathrm{H}}$, directly inserting (3.1) into (2.37) leads to
@@
For later convenience, we further define $\Omega_{\mathrm{GW}}^{\mathrm{H}}(k)=\sum_{i,j,l=1}^{n}\Omega_{\mathrm{GW}}^{\mathrm{H},ijl}(k)$ with
@@
To locate possible peaks in the non-Gaussian case, recall that in the Gaussian case, the locations of peaks coincide with the poles of $J(t,s,t,s)$. However, for the non-Gaussian case, contributions like $\Omega_{\mathrm{GW}}^{\mathrm{H}}$ admit an integration of $J(t,s,t,s)$. In this case, the poles of $J(t,s,t,s)$ do not always correspond to the peak locations. This is because that the integral over a logarithmic divergent pole can still be finite, for example, $\ln|x|$ has a logarithmic divergent pole at $x=0$, while $\int_{-a}^{a}\ln|x|\mathrm{d}x=2a(\ln a-1)$ is finite. As we will see later, it is more practical to find the domains of peaks (the ranges of $\tilde{k}$) instead of precisely locating the positions of peaks. To locate the domains of peaks, note that, when including the poles of $J(t,s,t,s)$ in the integration domain specified by those Heaviside step functions, the integration value can be significantly enhanced so as to include the potential peaks. Nevertheless, the poles of $J(t,s,t,s)$ are not guaranteed to be located within the integration domain. To include the poles of $J(t,s,t,s)$ in the integration domain, note that we have proved in the Appendix A the upper bound of the integration range in the $t$ direction is always smaller than the biggest zero point of $J(t,s,t,s)$ in the $t$ direction of its domain, and the value of $J(t,s,t,s)$ is always small except at the pole for $t\in[0,\text{the biggest zero point}]$, thus, the sufficient condition for having a peak is to include the logarithmic divergent pole in the domain of $t$ that is specified by those Heaviside step functions in the integrand, and hence we can obtain the domain of peaks as shown shortly below.
@@
This is very different from the Gaussian case, for example, when $m=1$, it roughly gives $n<7.46$, that is, when $\Omega_{\mathrm{GW}}^{\mathrm{G}}$ only has two peaks, $\Omega_{\mathrm{GW}}^{\mathrm{H}}$ can still have more than two peaks. In the left panel of Fig. 2, we present an example of induced GWs from primordial curvature perturbations with a double-$\delta$ peak. Taking $k_{*2}=4k_{*1}$, for instance, $\Omega_{\mathrm{GW}}^{\mathrm{G}}$ admits only two peaks while $\Omega_{\mathrm{GW}}^{\mathrm{H}}$ has at least three peaks (the other two peaks are overwhelmed, because their amplitudes are small compared to the other three peaks). In the right panel of Fig. 2, we let $m=1$ and vary $n$ from 7.5 to 6 to show that our sufficient condition (3.8) for the peak to appear is indeed fulfilled  [^6] In fact, this is a slightly weak condition, as it requires roughly $n<6$ to give a sizeable peak..
Due to the existence of integral, the peak of $\Omega_{\mathrm{GW}}^{\mathrm{H}}$ is usually broader than $\Omega_{\mathrm{GW}}^{\mathrm{G}}$ (like the plots in the left panel of Fig. 2), and hence it would be more practical to narrow down a tighter domain of peaks for $\Omega_{\mathrm{GW}}^{\mathrm{H}}$ rather than finding the precise locations of peaks as it is much more difficult. After some efforts, we can obtain such domains of peaks (see the Appendix A for details) as